What is a dual EM tensor?
1
Expert's answer
2019-06-10T09:43:31-0400

Dual electromagnetic tensor is defined as:


"F^{\mu \nu *} =\frac{1}{2}\epsilon^{\mu \nu\alpha \beta} F_{\alpha \beta}"

where "\epsilon^{\mu \nu\alpha \beta}" - Levi-Chivita tensor